Solve this application problem using a system of equations: In a basketball game, Alan scored 30 points, consisting only of three-point shots and two-point shots. He made a total of 11 shots. How many shots of each type did he make?
Answer by Theo(13342)

x = number of 2 point shots.
y = number of 3 point shots.
x + y = 11
2x + 3y = 30
multiply both sides of first equation by 2 and leave second equation alone to get:
2x + 2y = 22
2x + 3y = 30
subtract first equation from second to get:
y = 8
since x + y = 11, then x = 3
with x = 3 and y = 8, you get:
x + y = 3 + 8 = 11
2x + 3y = 6 + 24 = 30
solution is number of 2 point shots is 3 and number of 3 point shots is 8.